# Estimating Reaction Barriers with Deep Reinforcement Learning

### v1

@@ 3 Experiments
The proposed algorithm is applied to determine the pathway with the minimum energy barrier on the M"uller–Brown potential energy surface Müller and Brown [1979]. The M"uller–Brown potential has been used to benchmark the performance of several algorithms that determine the minimum energy pathways, such as the molecular growing string method Goodrow et al. [2009], Gaussian process regression for nudged elastic bands Koistinen et al. [2017], and accelerated molecular dynamics Vlachas et al. [2022]. Therefore, it is also used in this work to demonstrate the applicability of the proposed method. A custom Gym environment Towers et al. [2023] was created following the gymnasium interface (inheriting from the class Gym) to model the problem as a Markov Decision Process to be solved by a reinforcement learning pipeline. The values for the parameters used in Algorithm 1 are listed in Table 1.
@@ 3.1 Results
The M"uller–Brown potential is characterized by the following potential:
@@
where $W=(-200,-100,-170,15)$, $a=(-1,-1,-6.5,0.7)$, $b=(0,0,11,0.6)$, $c=(-10,-10,-6.5,0.7)$, $\overline{x}=(1,0,-0.5,-1)$, and $\overline{y}=(0,0.5,1.5,1)$. The potential energy surface for the system is plotted in Figure 3(a), and the local minima are at $(-0.558,1.442)$, $(0.623,0.028)$, and $(-0.050,0.467)$ with the value of $V(x,y)$ being $-146.7$, $-108.2$, and $-80.8$, respectively. The RL agent was trained to locate a path on this surface from $(0.623,0.028)$ with a random step (with zero mean and a standard deviation of $0.1$) added to it as the initial state to $(-0.558,1.442)$ as the terminal state, with the minimum energy barrier. The first random step was chosen to avoid the same starting point in each training iteration of the agent, so it learns a more generalized policy. Some of the parameters for the Markov Decision Process to model this potential are given in Table 2.
@@
(b) Learning curve with the shaded region representing one standard deviation from the mean over $11$ trials.
Figure 3: (Left) Environment in which the agent learns to find the path with the minimum energy barrier. (right) The rising learning curve against the number of evaluation episodes indicates that the agent learns to find the path with minimum energy barrier (recall that the reward was defined as the negative of the energy of a state, which leads to ascending learning curve and maximizing of the reward indicating that the agent learning a path with lower energy barrier). The agent was run in evaluation mode once every $10$ training steps.
(a) TD3
(b) SAC
(c) Algorithm 1
Figure 4: Scatter plot of the regions visited by the reinforcement learning agent during the course of learning while using different algorithms. While TD3 (left plot) does reach the intended target, the neural nets over-fits and take a more direct path to the terminal state and hence do not give the correct estimate of the minimum energy barrier for the transition. SAC (center plot) shows an improved performance and reaches the intended target. However, while generating trajectories in the testing environment, most of the trajectories did not leave the local minima in the vicinity of the start state. Moreover, the learnt policy has a high variance. The proposed Algorithm 1 learns a much stable policy and confines itself to exploring the region with lower energies leading to the terminal state rather than the entire environment. It explores sufficiently and then exploits the state-action values learnt appropriately giving better estimates of the energy barrier for the transition.
@@
Table 2: Some parameters of the Markov Decision process to find pathways with the minimum energy barrier on the chosen potential.
(a) Trajectories generated by the agent after training.
(b) Energy profiles for the generated trajectories.
Figure 5: Trajectories generated by the trained agent following the learnt policy.
@@
In Figure 5(a), an ensemble of paths generated by the trained RL agent with the starting points slightly perturbed from $(0.623,0.028)$ by noise added from $\mathcal{N}(0,0.1)$ is plotted on the energy surface. The energy profiles along the generated trajectories are plotted in Figure 5(b) aligned by the maximum of the profiles (and not by the start of the trajectories) for better visualization. The predicted energy barrier for the transition of interest is $-40.36\pm 0.21$. One can see that the agent learns to predict the path with the correct minimum energy barrier, albeit the energy barrier estimated by the agent is a little higher than the optimal analytical solution ($-40.665$). However, the result demonstrates that reinforcement learning algorithms can be used to locate the minimum energy barrier for transitions between stable states in complex systems. The paths suggested by the trained agent cluster around the minimum energy path and pass through the vicinity of the actual saddle point representing the energy barrier. However, there still seems to be some way to go to improve the sampling densities around the saddle point, which determines the barrier height, to avoid overestimating it.
@@ 3.2 Ablation Studies
Several modifications were made to the standard SAC algorithm to be used in this particular case (highlighted in blue in Algorithm 1). Studies were performed to understand the contribution of each individual component to the working of the algorithm in this particular environment by comparing the performance of the algorithm with different hyperparameters for a component. Each modification and its contribution to the overall learning of the agent are described in the following sections. The mean and the standard deviation of the returns from the last $100$ training steps for each modification to the existing algorithm are listed in Table 6(a) to compare the performance of the agents. The modification that leads to the highest returns is highlighted.
@@
[…]
00 $28534\pm 8137$
$2$ $55438\pm 1158$
@@ 3.2.3 Tuning the Entropy Coefficient
The entropy coefficient $\alpha$ can be tuned as the agent learns (as done in line 18 of Algorithm 1), which overcomes the problem of finding the optimal value for the hyperparameter $\alpha$ Haarnoja et al. [2019]. Moreover, simply fixing $\alpha$ to a single value might lead to a poor solution because the agent learns a policy over time: it should still explore regions where it has not learned the optimal action, but the policy should not change much in regions already explored by the agent that have higher returns. In Figure 6(d), the effect of the variation of the hyperparameter $\alpha$ on the learning of the agent is compared. As can be seen, a tunable $\alpha$ allows the agent to learn steadily, encouraging it to explore more in the earlier episodes and exploiting the returns from these explored regions in the latter episodes, resulting in a more stable learning curve (blue curve). A too low value of $\alpha$, such as $10^{-3}$ or $10^{-1}$, makes the algorithm more deterministic (TD3-like), which leads to sub-optimal performance and the agent being stuck in a local minimum (plotted as green and red curves, respectively). An $\alpha$ value of $0.1$ has comparable performance to the tunable $\alpha$, but the learning curve is less stable and there are abrupt changes in the policy function (magenta curve). The original implementation of SAC suggested $0.2$ as a fixed value for $\alpha$, which leads to a learning curve resulting in a policy with high variance (yellow curve). A too high value of $\alpha$, such as $0.5$, makes the algorithm more stochastic (REINFORCE-like), which also leads to sub-optimal learning (cyan curve).
4 Discussions and Conclusion
@@
Figure 7: (Left) Three possible pathways on a potential energy surface (taken from E et al. [2007]) passing though different saddle points. A simple interpolation between the starting and ending points would lead to the dotted trajectory. Running traditional gradient based methods would not improve this pathway as it passes through a local minima. This pathway has an energy barrier of $1.50$. The pathway with the lowest energy barrier, $1.27$, is depicted by a solid line, and was learnt by the reinforcement learning agent following Algorithm 1. (right) The energy profile for the three possible transition pathways. Nudged elastic band would have overestimated the energy barrier for the transition by $(150-127)/127$ or $18\%$, and hence underestimates the frequency with which it occurs by $1-e^{-1.50-(-1.27)}=20\%$.
@@
Advancements in reinforcement learning algorithms based on the state-action value function have led to their application in diverse sequential control tasks such as Atari games, autonomous driving, and robot movement control. This project formulated the problem of finding the minimum energy barrier for a transition between two local minima as a cost minimization problem, solved using a reinforcement learning setup with neural networks as function approximators for the actor and critics. A stochastic policy was employed to facilitate exploration by the agent, further perturbed by random noise. Target networks, delayed updates of the actor, and a replay buffer were used to stabilize the learning process for the reinforcement learning agent. While the proposed framework samples the region around the saddle point sufficiently, providing a good estimate of the energy barrier for the transition, there is definitely scope for improvement. As future work, the method could be applied to more realistic systems.
Previous works in determining transition pathways using deep learning or reinforcement learning techniques include formulating the problem as a shooting game solved using deep reinforcement learning Zhang et al. [2021]. Additionally, in Holdijk et al. [2023], this problem is formulated as a stochastic optimal control problem, where neural network policies learn a controlled and optimized stochastic process to sample the transition pathway using machine learning techniques. Stochastic diffusion models have been used to model elementary reactions and generate the structure of the transition state, preserving the required physical symmetries in the process Duan et al. [2023]. Furthermore, the problem of finding transition pathways was recast into a finite-time horizon control optimization problem using the variational principle and solved using reinforcement learning in Guo et al. [2024]. Recent work Barrett and Westermayr [2024] used an actor-critic reinforcement learning framework to optimize molecular structures and calculate minimum energy pathways for two reactions.
@@
This work differs from previous efforts by providing a much simpler formulation of the problem, using the energy of the state directly as the reward while searching for transition pathways with the minimum energy barrier. One of the main advantages of this method is that, unlike traditional methods such as the nudged elastic band or the growing string method, it does not require an initial guess for the transition pathway. Traditional methods use energy gradient information along the pathway to iteratively improve to a pathway with better energetics. However, the success of these methods depends on the initial guess for the pathway, which might be stuck in a local minimum, as shown in Figure 7, leading to a sub-optimal solution. This could result in an overestimate of the energy barrier and subsequently an underestimate of the probability for the transition to occur, leading to imperfect modeling of the system. On the other hand, the use of a stochastic policy in a reinforcement learning setup avoids this problem, increasing the chances of finding a better estimate of the transition barrier as the agent explores the state space. However, as a trade-off for the simple model and generic approach, the agent learns slowly and requires a large number of environment interactions.

### v2

@@ 3 Experiments
The proposed algorithm is applied to determine the pathway with the minimum energy barrier on the Müller–Brown potential energy surface Müller and Brown [1979]. The Müller–Brown potential has been used to benchmark the performance of several algorithms that determine the minimum energy pathways, such as the molecular growing string method Goodrow et al. [2009], Gaussian process regression for nudged elastic bands Koistinen et al. [2017], and accelerated molecular dynamics Vlachas et al. [2022]. Therefore, it is also used in this work to demonstrate the applicability of the proposed method. A custom Gym environment Towers et al. [2023] was created following the gymnasium interface (inheriting from the class Gym) to model the problem as a Markov Decision Process to be solved by a reinforcement learning pipeline. The values for the parameters used in Algorithm 1 are listed in Table 1.
@@ 3.1 Results
The Müller–Brown potential is characterized by the following potential:
@@
where $W=(-200,-100,-170,15)$, $a=(-1,-1,-6.5,0.7)$, $b=(0,0,11,0.6)$, $c=(-10,-10,-6.5,0.7)$, $\overline{x}=(1,0,-0.5,-1)$, and $\overline{y}=(0,0.5,1.5,1)$. The potential energy surface for the system is plotted in Figure 3(a), and the coordinates of the local minima and saddle points for the potential energy surface and their corresponding energies are tabulated in Table 3(c). The RL agent was trained to locate a path on this surface from $S\,(0.623,0.028)$ with an initial random step (with zero mean and a standard deviation of $0.1$) taken as the starting state to $T\,(-0.558,1.442)$ as the terminal state, with the minimum energy barrier. The first random step was chosen to avoid the same starting point in each training iteration of the agent, so it learns a more generalized policy. Some of the parameters for the Markov Decision Process to model this potential are given in Table 3(c).
@@
pt x y E
S $0.623$ $0.028$ $-108.2$
Minima T $-0.558$ $1.442$ $-146.7$
U $-0.050$ $0.467$ $-80.8$
Saddle V $0.210$ $0.300$ $-72.3$
points W $-0.820$ $0.620$ $-40.7$
(b) Minima and saddle points on the chosen potential energy surface.
@@
(c) Some parameters of the Markov Decision process to find pathways with the minimum energy barrier on the chosen potential.
Figure 3: The environment in which the agent learns to find the path with the minimum energy barrier.
(a) TD3
(b) SAC
(c) Algorithm 1
Figure 4: Scatter plot of the regions visited by the reinforcement learning agent during the course of learning while using different algorithms.
(a) Learning curve for the agent averages over $11$ trials.
(b) Least reward collected in the episode.
(c) Trajectories generated by the agent after training.
(d) Energy profiles for the generated trajectories.
Figure 5: (a) The learning curve for the agent in the reinforcement learning environment. (b) The plot of the variation of the least reward collect by the agent in a step with the validation episode count. (c) Trajectories generated by the trained agent following the learnt policy along with the corresponding energy profiles (d).
3.1.1 Comparison of different algorithms
Figure 4 shows scatter plots of the trajectories generated by various reinforcement learning algorithms: TD3 in Figure 4(a), SAC in Figure 4(b), and the proposed modified SAC algorithm in Figure 4(c). While the agent trained by the TD3 algorithm does reach the intended target state, it starts to exploit a flaw in the formulation of the MDP by trying to reach the vicinity of the final state quickly and staying near enough to it so that it collects rewards, but does not terminate the episode. This results in a high density in the plot along the straight line connecting the initial and final states and around the final state. It gives a much higher estimate than the correct minimum energy barrier for the transition. The agent trained using SAC shows improved performance, possibly due to the entropy regularizer forcing it to learn a more diverse policy (rather than one that would result in a straight line connecting the initial and final states). However, while generating trajectories in the testing environment, most of the trajectories did not leave the local minima in the vicinity of the start state. Moreover, the learned policy has high variance. The proposed Algorithm 1 learns a much more stable policy and confines itself to exploring the region with lower energies leading to the terminal state (specifically the vicinity of the saddle point) rather than the entire environment. It explores sufficiently and then exploits the state-action values learned appropriately, providing better estimates of the energy barrier for the transition.
The learning curve for the agent under Algorithm 1 is shown in Figure 5(a). The data for this curve were generated by allowing the agent to solve the MDP in evaluation mode once every $10$ training episodes, where the neural networks were not updated to monitor the agent’s learning. The ascending learning curve indicates that the agent gradually learns to find a path to the terminal state that maximizes the rewards. The blue line represents the median reward, while the green line shows the mean reward over $11$ training iterations, each consisting of $10\times 1000$ training episodes. The light blue shaded region denotes the spread of the rewards (maximum and minimum). A low spread in rewards indicates consistent performance by the reinforcement learning agent in the validation episodes.
3.1.2 Performance of the trained agent
In Figure 5(c), an ensemble of paths generated by the trained RL agent is plotted on the surface of the potential energy, with the starting points slightly perturbed from $(0.623,0.028)$ by noise added from $\mathcal{N}(0,0.1)$ on the surface of the potential energy. The model used to generate these trajectories was the model at the $500$th validation step (and not after the entire training consisting of $1000$ validation steps) for reasons elaborated later. It can be seen from Figure 5(c) that the agent spends more time near the terminal state rather than reaching the terminal state to receive an immediate reward, as it allows the agent to collect rewards for more steps. In the case of a coarse-grained maze representation of the potential energy surface, this problem was solved by rewarding the agent only the first time it visited a grid cell. Using a similar idea of not rewarding the agent when it is in the close neighborhood of an already visited state artificially perturbs the reward map and did not work in this case. The best performance was achieved by gradually varying the maximum number of steps the agent was allowed to take in an episode. If the number of steps allowed in an episode is too short, the agent does not escape the local minima to reach the terminal state. If the number of steps is too large, then the agent reaches the terminal state and discovers that it receives larger rewards by remaining in its vicinity, but not so close that the episode is terminated. In an attempt to reach the terminal state earlier, the agent tries to approach the terminal state sooner, choosing a more direct pathway, which lifts the trajectory out of the saddle point. It was observed that a maximum of $500$ steps per episode led to the best performance by the agent. The agent did not leave the local minima if fewer steps were allowed ($200$ steps), and the agent passed through states with much higher energies than optimal to reach a minimum energy state if longer episodes were allowed ($1000$ steps).
@@ 3.1.2 Performance of the trained agent
Early stopping during the training of a neural network has often been found to be helpful in scenarios where continued training worsens the performance of the model Nakkiran et al. [2020], Bai et al. [2021]. Borrowing the idea of early stopping, the agent’s training was stopped when the minimum reward collected by the agent (corresponding to the maximum energy along the pathway) started increasing again. This behavior was observed in the case of the agent, as plotted in Figure 5(b). The minimum reward collected by the agent during the episode increases initially (indicating that the agent finds a pathway with a progressively better energy barrier) until the $500$th validation step before decreasing slightly. Plots for only $4$ of the $11$ trials are shown for clarity. This might indicate that the agent does not improve its performance after that step. Furthermore, the learning curve in Figure 5(a) shows an increase in spread after $500$ iterations. These reasons led to using the model after $500$ iterations to generate the final trajectory in test mode to estimate the energy barrier for the reaction. The energy profiles along the generated trajectories are plotted in Figure 5(d) aligned by the maximum of the profiles (and not by the start of the trajectories) for better visualization. The energy barrier predicted for the transition of interest is $-40.36\pm 0.21$. One can see that the agent learns to predict the path with the correct minimum energy barrier, albeit the energy barrier estimated by the agent is a little higher than the optimal analytical solution ($-40.665$). However, the result demonstrates that reinforcement learning algorithms can be used to locate the minimum energy barrier for transitions between stable states in complex systems. The paths suggested by the trained agent cluster around the minimum energy path and pass through the vicinity of the actual saddle point that represents the energy barrier. However, there still seems to be some way to go to improve the sampling densities around the saddle point, which determines the barrier height, to avoid overestimating it.
3.1.3 On the choice of the scaling factor
The scaling factor, $\lambda$, scales the action for the agent. In most cases, it is used to adjust the step size for the agent while keeping the action space within some standard interval ($[0,1]$ or $[-1,1]$). This scaling factor was varied along with the number of steps in an episode, and the combination of $0.01$ for the $\lambda$ and $500$ steps in an episode led to the best performance of the agent. With these parameters, the agent reaches near the terminal state with just the number of required small enough steps to end the episode. A larger value of $\lambda$ resulted in the agent taking longer steps over regions of the potential energy surface with a higher energy to give an incorrect estimate of the barrier height. Smaller values of $\lambda$ led to smaller steps, and the agent did not leave the local minima to explore other regions of the potential energy surface, and is unsuccessful a its assigned task.
As noted at the end of Section 2.1, the formulated Markov decision process suffers from the drawback that that agent might stay at a small distance $\epsilon(>\delta)$ from that target state, and collect rewards until the remainder of the episode. To discourage the agent from doing this, the episode is truncated after $500$ steps. Increasing the number of steps in the episode would encourage the agent to stay a small distance away from the target state, rather than reaching the target state and terminating the episode, once it realizes that it can increase the total reward collected by this. On the other hand, if the number of steps in an episode is too small, it would not reach near the target state. The choice of the scaling factor $\lambda$ is related to the number of steps in an episode. Together, they determine the maximum distance from the start state the agent can reach. If both $\lambda$ and the number of steps in an episode is decreased below a limit, the agent would never be able to reach the target state  [^1] For example, if the maximum number of steps for the maze in Figure 2 is limited to $5$, then the agent can never reach the terminal state. As an analogy, decreasing the cell size, resulting in increasing the number of cells in the maze, would be equivalent to decreasing $\lambda$ in the current case.. With an appropriate choice of $\lambda$ and the length of the episode, the agent does not make too many long jumps through higher energy states to reach a state with lower energy faster. Decreasing $\lambda$ would require increasing the maximum number of steps in an episode, so that the agent explores regions away from the starting point, but not too much so that the trajectory passes through regions with higher energy. A few experiments were done to determine the pair of the values for $\lambda$ and the number of steps in an episode which gives the best performance by the agent, and the results are plotted in Appendix A.
@@ 3.2 Ablation Studies
Several modifications were made to the standard SAC algorithm to be used in this particular case (highlighted in blue in Algorithm 1). Studies were performed to understand the contribution of each individual component to the working of the algorithm in this particular environment by comparing the performance of the algorithm with different hyperparameters for a component. The parameters for one modification were varied, keeping the parameters for the other two modifications unchanged from the fine-tuned algorithm. Each modification and its contribution to the overall learning of the agent are described in the following sections. The mean and the standard deviation of the returns from the last $100$ training steps for each modification of the existing algorithm are listed in Table 6(a) to compare the performance of the agents. The modification that leads to the highest returns is highlighted.
@@
smoothing present $\mathbf{8651\pm 107}$
00 $4826\pm 984$
$2$ $7738\pm 234$
policy update $4$ $7741\pm 220$
delay $\mathbf{8}$ $\mathbf{7994\pm 206}$
$16$ $7677\pm 204$
$32$ $5783\pm 142$
@@ 3.2.3 Tuning the Entropy Coefficient
The entropy coefficient $\alpha$ can be tuned as the agent learns (as done in line 18 of Algorithm 1), which overcomes the problem of finding the optimal value for the hyperparameter $\alpha$ Haarnoja et al. [2019]. Moreover, simply fixing $\alpha$ to a single value might lead to a poor solution because the agent learns a policy over time: it should still explore regions where it has not learned the optimal action, but the policy should not change much in regions already explored by the agent that have higher returns. In Figure 6(d), the effect of the variation of the hyperparameter $\alpha$ on the learning of the agent is compared. As can be seen, a tunable $\alpha$ allows the agent to learn steadily, encouraging it to explore more in the earlier episodes and exploiting the returns from these explored regions in the latter episodes, resulting in a more stable learning curve (blue curve). A too low value of $\alpha$, such as $10^{-3}$ or $10^{-2}$, makes the algorithm more deterministic (TD3-like), which leads to suboptimal performance and the agent being stuck in a local minimum (plotted as green and red curves, respectively). An $\alpha$ value of $0.1$ has a performance comparable to the tunable $\alpha$, but the learning curve is less stable and there are abrupt changes in the policy function (magenta curve). The original implementation of SAC suggested $0.2$ as a fixed value for $\alpha$, which leads to a learning curve that results in a policy with high variance (yellow curve). A too high value of $\alpha$, such as $0.5$, makes the algorithm more stochastic (REINFORCE-like), which also leads to suboptimal learning (cyan curve).
3.3 Some more surfaces
Here we demonstrate the results by an agent trained by Algorithm 1 on some more two-dimensional potential energy surfaces with two potential wells.
$V(x,y)$ $=(x-1)^{4}+(y-1)^{4}+(x+1)^{4}+(y+1)^{4}-20(x-y)^{2}+60x$ (2)
$V(x,y)$ $=\left[y-0.4(x^{2}-4x)\right]^{2}-x^{2}+0.1(x^{4}+y^{4})+0.5x$ (3)
$V(x,y)$ $=\left[(1-x^{2}-y^{2})^{2}+\frac{y^{2}}{x^{2}+y^{2}}\right]\left(1+\frac{1}{1+e^{-y}}\right)$ (4)
While the potential energy surfaces represented by Equation 4 was taken from Vanden-Eijnden et al. [2007], those represented by Equation 2 and Equation 3 were crafted by hand. The results of using the proposed Algorithm 1 on these potential energy surfaces are depicted in Figure 7. The left column shows the trajectories generated by the agent after training between the two minima on the potential energy surface. The central column plots the learning curves of the agent, with the mean cumulative sum of returns plotted as a green line. The spread of the sum of the returns is shaded in blue, while the rolling mean is represented as a solid blue line. The right column shows the energy profiles of the generated trajectories. The estimated energy barrier for the transition on the potential energy surface given by Equation 2 was $-5.575$ while that for the potential energy surface given by Equation 3 was $0.94$.
(a) Equation 2
(b) Equation 3
Figure 7: Results on using the proposed algorithm on the potential energy surfaces depicted by (a) Equation 2 and (b) Equation 3.
4 Discussions
(Horizon Europe Research and Innovation Project CORENET) Previous work in determining transition pathways using deep learning or reinforcement learning techniques includes formulating the problem as a shooting game solved using deep reinforcement learning Zhang et al. [2021]. The authors in Zhang et al. [2021] sample higher energy configurations and shoot trajectories with randomized initial momenta in opposite directions, expecting them to converge at the two desired local minima. In contrast, the method proposed here starts from a minimum on the potential energy surface and attempts to generate a trajectory to another minimum. Additionally, in Holdijk et al. [2023], the problem is formulated as a stochastic optimal control problem, where neural network policies learn a controlled and optimized stochastic process to sample the transition pathway using machine learning techniques. Stochastic diffusion models have also been used to model elementary reactions and generate the structure of the transition state, preserving the required physical symmetries in the process Duan et al. [2023]. Furthermore, the problem of finding transition pathways was recast into a finite-time horizon control optimization problem using the variational principle and solved using reinforcement learning in Guo et al. [2024]. Moreover, a hybrid-DDPG algorithm was implemented in Mills et al. [2022] to identify the global minimum on the Müller–Brown potential, but did not identify pathways between minima as in this work. Recent work Barrett and Westermayr [2024] used an actor-critic reinforcement learning framework to optimize molecular structures and calculate minimum energy pathways for two reactions.
There has also been previous work Zhou et al. [2017] to optimize chemical reactions by perturbing the experimental conditions to achieve better selectivity, purity, or cost for the reaction using deep reinforcement learning. While this approach has macroscopic applications in laboratory settings, the method proposed here focuses on a much narrower problem: given a potential energy surface, how well can the minimum energy barrier be estimated for a transition between two minima? Deep reinforcement learning has also been used to find a minimum energy pathway consisting of multiple elementary transitions in catalytic reaction networks Lan and An [2021]. While the free energy barrier for a transition (which is mapped to a reward) between two states is calculated using density functional theory (DFT) with VASP software in Lan and An [2021], Lan et al. [2024], the objective of the proposed method in this work is to estimate that free energy barrier using an agent trained via deep reinforcement learning, which does not require quantum mechanical calculations. In addition, reinforcement learning techniques are implemented in Zhang and Lapkin [2023] to minimize the cost of synthesis pathways (consisting of multiple elementary transitions) considering the price of the starting molecules and the atom economy of individual transitions. Furthermore, a reinforcement learning approach is used to search for process routes that optimize economic profit for a Markov decision process that models the thermodynamic state space as a graph.
5 Conclusion
@@
Figure 8: (a) Three possible pathways on a potential energy surface given by Equation 4 passing though different saddle points and (b) the energy profile for the three possible transition pathways.
@@
Advancements in reinforcement learning algorithms based on the state-action value function have led to their application in diverse sequential control tasks such as Atari games, autonomous driving, robot movement control, and more physical domains Wen et al. [2023], Gow et al. [2022], Beeler et al. [2023, 2021]. This project formulated the problem of finding the minimum energy barrier for a transition between two local minima as a cost minimization problem, solved using a reinforcement learning setup with neural networks as function approximators for the actor and critics. A stochastic policy was employed to facilitate the exploration by the agent, further perturbed by random noise. Target networks, delayed updates of the actor, and a replay buffer were used to stabilize the learning process for the reinforcement learning agent. While the proposed framework samples the region around the saddle point sufficiently, providing a good estimate of the energy barrier for the transition, there is definitely scope for improvement. The method has been applied only to a two-dimensional system, but as a future work, it could be extended to more realistic and higher-dimensional systems. One promising alternative would be to use max-reward reinforcement learning Veviurko et al. [2024], as it aligns well with the objective of maximizing the minimum reward obtained in an episode. However, a drawback of this method is that the reinforcement learning agent must be trained from scratch if one needs to find minimum energy pathways on a different potential energy surface. In other words, an agent trained on one potential energy surface cannot be used to determine the minimum energy barrier on a different surface, similar to how an agent trained in one Gymnasium environment cannot solve tasks in another. Another limitation is that the agent can only locate pathways to minima lower than the starting minimum; otherwise, remaining at the starting position minimum would yield higher rewards for the agent.
@@
This work differs from previous work that uses reinforcement learning Zhang et al. [2021], Barrett and Westermayr [2024], Guo et al. [2024], Holdijk et al. [2023] by providing a much simpler formulation of the problem, using the energy of the state directly as the reward while searching for transition pathways with the minimum energy barrier. One of the main advantages of using a reinforcement learning based method is that, unlike traditional methods such as the nudged elastic band or the growing string method, it does not require an initial guess for the trajectory. Traditional methods use energy gradient information along the trajectory to iteratively improve to a trajectory with better energetics. However, the success of these methods depends on the initial guess for the trajectory, and gradient-based methods might get stuck in a local minimum. As shown on the potential energy surface represented by Equation 4 in Figure 8, there may be multiple saddle points between two minima. The trajectory to which a nudged elastic band or a growing string method converges depends on the initial guess of the starting trajectory. Typically, the initial guess trajectory is a simple linear interpolation between the starting and ending points, which leads to the dotted trajectory (path $2$ with a barrier of $1.50$ units). Traditional gradient-based methods report this trajectory as the optimal one because the local gradients along the trajectory are minimal and cannot be improved by perturbation. However, the reinforcement learning-based method proposed in this work identifies the trajectory represented by a solid line (Path $3$ with a barrier of $1.27$ units) as the minimum energy pathway. The suboptimal solution overestimates the energy barrier for the transition by $(150-127)/127$ or $18\%$, and therefore underestimates the frequency with which it occurs by $1-e^{-1.50-(-1.27)}=20\%$. Underestimates of the probability for a transition to occur would lead to imperfect modeling of the dynamics of the system. The use of a stochastic policy in a reinforcement learning setup avoids this problem, increasing the chances of finding a better estimate of the transition barrier as the agent explores the state space. However, as a trade-off for the simple model and generic approach, the agent learns slowly, requires a large number of environment interactions, and would have to be retrained to work in a new potential energy surface.
Acknowledgments and Disclosure of Funding
The author would like to gratefully acknowledge the computational resources provided by Institut for Matematik og Datalogi, Syddansk Universitet for this work. The work was also supported by the European Union (Horizon Europe Research and Innovation Project CORENET) and the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract numbers 22.00017 and 22.00034, respectively. The author also thanks the two reviewers for their insightful suggestions to improve the manuscript.
Appendix A Additional heatmap plots
A few experiments by varying the scaling factor for the action, $\lambda$, and the number of steps in an episode, $n$, were performed. The regions of the potential energy surface explored by the agent under those conditions are plotted in Figure 9. With a small value for $\lambda$, the agent does not climb out of the local minima containing the initial state (Figure 9(a)), while with a large value for $\lambda$, the agent jumps over high energy regions of the potential energy surface in the bid to reach a low energy state faster (Figure 9(f)), giving an incorrect estimate of the energy barrier for the transition.
(a) $\lambda=0.005,n=500$
(b) $\lambda=0.005,n=1000$
(c) $\lambda=0.005,n=100$
(d) $\lambda=0.001,n=200$
(e) $\lambda=0.02,n=100$
(f) $\lambda=0.02,n=500$
Figure 9: Comparative scatter plots of the regions visited by the agent with different values for $\lambda$ and $n$.